The fishing pole makes an angle of 20.0° with the horizontal. What is the torque exerted by the fish about an axis perpendicular to the page and passing through the angler's hand if the fish pulls on the fishing line with a vector force F=100 N at an angle 37.0° below the horizontal? The force is applied at a point 2.00 m from the angler's hands.


1
Expert's answer
2020-11-10T06:59:09-0500

First, calculate the angle between the force and the fishing pole:



As we see, the angle is


"\theta=37°+20°=57°."

The torque is caused by the component of force perpendicular to the pole:


"\tau=F\text{ sin}\theta\cdot L=100\text{ sin}57°\cdot2=168\text{ Nm}."